# Kolmogorov-Arnold Energy Models: Fast and Interpretable Generative Modeling

### v6

@@ 1 Introduction
Deep generative modeling is a core subfield of machine learning that enables systems to learn complex data distributions, generate novel data, and support a wide range of tasks including simulation, compression, and representation learning. Two prominent and widely studied approaches are Variational Autoencoders (VAEs) (Kingma & Welling, 2022) and Generative Adversarial Networks (GANs) (Goodfellow et al., 2014). Both frameworks assume that each data point is generated from a low-dimensional latent vector via a top-down neural network, where the latent vector is typically drawn from a simple, non-informative prior, such as a uniform or isotropic Gaussian distribution.
@@
Latent energy-based prior models (EBMs) (Pang et al., 2020) extend this framework by learning an expressive, data-dependent prior directly in the latent space. This prior can be trained jointly with the top-down generator using maximum likelihood estimation (MLE), where the EBM learns by the statistical discrepancy between samples from the prior and posterior distributions, following a paradigm similar to Empirical Bayes. Each learning iteration requires Markov Chain Monte Carlo (MCMC) sampling from both the prior and posterior, specifically the Unadjusted Langevin Algorithm (ULA) (Brooks et al., 2011; Rossky et al., 1978b; Roberts & Stramer, 2002), which offers greater modeling flexibility compared to the amortized inference methods of VAEs.
@@
Despite being extendable to several high-impact research directions, such as multimodal generative modeling (Yuan et al., 2024), and interpretable classification and meta-learning (Kong et al., 2021), (which involves rapid adaptability and few-shot generalization to out-of-distribution tasks), latent EBMs face critical limitations due to their reliance on MCMC sampling.
@@
Firstly, MCMC is inherently iterative, and algorithms such as ULA require repeated gradient passes through the model, which introduces significant computational overhead during both training and inference. Secondly, ULA is best suited for sampling from smooth, unimodal distributions and can struggle to explore the multimodal latent distributions that are typical in latent variable generative modeling. Moreover, the framework presented by Pang et al. (2020) requires separate ULA samplers for the prior and posterior. This introduces additional experimental complexity since ULA is sensitive to hyperparameter tuning: step sizes that are too small lead to poor mixing and convergence, while larger step sizes can accumulate discretization bias due to its use of the Euler-Maruyama method (Wilkie, 2004) to evolve the underlying Langevin equation.
@@
Additionally, it is not evident how the unique separation between the prior and generator networks can be leveraged for interpretability or training efficiency. Unlike the framework introduced by Karniadakis et al. (2021), which describes how scientific inductive and learning biases can be embedded into neural networks to reduce their reliance on data and improve training efficiency, generalization, and explainability, (all of which are critical for ensuring trust in Artificial Intelligence), it remains unclear how the EBM prior can be effectively aligned with the posterior to achieve similar effects. Although ideas from mechanistic interpretability, as reviewed by Bereska & Gavves (2024), could potentially be applied, there is currently no established method to leverage the unique strengths of the EBM prior framework for this purpose.
@@ 1.1 Inductive biases and fast sampling
To address these limitations, we reimagine the EBM prior into a novel univariate or element-wise model, drawing inspiration from Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024) and deep spline networks Bohra et al. (2020), which replace the sigmoidal activations used in multilayer perceptrons Murtagh (1991) with spline interpolations. By representing the latent space with univariate functions and distributions, we enable fast, (single forward pass), and exact prior sampling with the inverse transform method (Devroye, 1986) and improve interpretability for domain experts who may not be deeply familiar with computational statistics and machine learning. This allows the EBM prior to be aligned with the posterior before training begins.
In Sec. 3.1, we show that adequately chosen priors can enable efficient training with importance sampling (IS) (Goertzel, 1949). This is a technique for estimating posterior expectations that is otherwise avoided in the literature due to its high variance under severe prior-posterior mismatch and the curse of dimensionality, (Dupuis & Wang, 2004), which can be avoided in our model due to the low dimensionality of the latent space. Moreover, we show that the trained priors are easy to visualize and recover, offering new opportunities for scientific discovery and knowledge transfer to models with different latent structures.
Multiple KAN bases and reference priors were used to verify the effect of inductive bias on training with IS. However, Gaussian Radial Basis Functions (RBF KANs) (Li, 2024), were adopted instead of the cubic B-spline bases proposed by Liu et al. (2024), as RBFs are more efficient for GPU implementation, and Poggio & Girosi (1990) previously demonstrated that they provide an optimal solution to the regularized approximation/interpolation problem.
To derive our new model, which we call the Kolmogorov-Arnold Energy Model (KAEM), we first introduce a novel adaptation of the Kolmogorov-Arnold Representation Theorem (KART) (Givental et al., 2009) by reinterpreting its inner functions as Markov kernels between probability spaces. Given its finite and definitive nature, KART can provide structural bias and reduce the reliance of architecture design on intuition and automated tuning.
@@ 1.2 Population-based Langevin sampling and Thermodynamic training
However, while ITS enables fast inference, posterior sampling must still be conducted using ULA when the IS estimator is insufficient. Prior work addressing poor MCMC convergence and multimodal exploration in EBMs has typically relied on diffusion models, such as the hierarchical EBM by Cui & Han (2024) and the diffusion-assisted EBM by Zhang et al. (2023). To maintain prior interpretability and fast inference post-training, we instead introduce posterior annealing as a novel and alternative strategy. This involves decomposing the posterior into a sequence of power posteriors, presented by Friel & Pettitt (2008), and recasting ULA with population-based MCMC methods (Swendsen & Wang, 1986; Marinari & Parisi, 1992; Hukushima & Nemoto, 1996). can The training criterion is derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by Calderhead & Girolami (2009); Annis et al. (2019) to ensure the impact of annealing on learning dynamics and parameter updates. Although this might be less expressive than hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training; and unlike denoising, which scales sequentially, annealing can scale by adding more temperature in parallel.
While ULA is computationally efficient and enables more flexible chain dynamics, it does not preserve detailed balance or reversibility. This can be recovered by incorporating Metropolis-Hastings (MH) adjustments (Metropolis & Ulam, 1949) via the Metropolis-adjusted Langevin Algorithm (MALA) (Rossky et al., 1978a). The code in Sec. A.1 optionally supports MALA with locally adaptive step-size tuning, (autoMALA) (Biron-Lattes et al., 2024), to assist when a single step size fails to accommodate the varying geometries of power posteriors, and when the MH criterion deteriorates under the curse of dimensionality, (Hairer et al., 2014). This is discussed in Secs. A.6.3 and A.6.4.
However, ULA remains the preferred method in our work, motivated by its efficiency and effectiveness, demonstrated by Pang et al. (2020) and further supported by the findings of Nijkamp et al. (2019b; a), with regards to how short-run, biased MCMC methods can still be highly effective for sampling from EBMs.
@@
The experiments in Sec. 3.2 demonstrate that our mixture prior model is able to generate coherent, (although low-quality), samples after a few training epochs on RGB images, while still supporting efficient, exact sampling via component-wise ITS. In contrast, annealing incurred substantial computational overhead, but produced little to no improvements in sample quality, rendering its use difficult to justify. However, this result may change as a larger number of power posteriors, longer training durations, and more advanced posterior sampling algorithms become practical with the hardware advances proposed by Zetta (2024). Their architecture enables fine-grained control over data flow and parallelized element-wise nonlinearity which is naturally well suited to supporting KAEM at scale, while improving the potential of univariate priors with its function fusion features.
@@ 2 The Kolmogorov-Arnold Energy Model
The Kolmogorov-Arnold Representation theorem (KART) (Givental et al., 2009) forms the basis of KAEM. A structured proof of the theorem has been provided by Dzhenzher & Skopenkov (2022).
@@ 2.1 Univariate energy-based prior
We interpret $u_{p}\sim\mathcal{U}(u_{p};\,0,1)$ in Eq. 1 as a uniform random variable in the probability space, $\left([0,1],\mathcal{B}([0,1]),\mu\right)$, where $\mathcal{B}\left([0,1]\right)$ is the Borel $\sigma$-algebra on $[0,1]$ and $\mu$ is the Lebesgue measure on $[0,1]$. To ensure that the transformation $u_{p}\mapsto\psi_{q,p}(u_{p})$ is well defined in a probabilistic setting, we require it to be a Markov kernel. For all Borel measurable sets $A\subseteq\mathcal{Z}$, where $\mathcal{Z}\subset\mathbb{R}$ is introduced as the latent space of KAEM, the map
$u_{p}\ \mapsto\ K(u_{p},A)=\delta_{\psi_{q,p}(u_{p})}(A)$ (2)
is measurable. Here, $\delta_{\psi_{q,p}(u_{p})}$ denotes the Dirac measure centred at $\psi_{q,p}(u_{p})$, ensuring the kernel is deterministic: i.e., once $u_{p}$ is sampled, the outcome in $\mathcal{Z}$ is fixed. This ensures the pushforward,
$\left([0,1],\mathcal{B}\left([0,1]\right),\mu\right)\;\to\;\left(\mathcal{Z},\mathcal{B}\left(\mathcal{Z}\right),\pi\right),$ (3)
is well defined and that no additional stochasticity is introduced beyond the uniform source variable, $u_{p}$. Here, $\mathcal{B}(\mathcal{Z})$ is the Borel $\sigma$-algebra on the latent space, $\mathcal{Z}$, and $\pi$ is the latent measure. The functions $\psi_{q,p}$ can therefore be conveniently interpreted as inverse cumulative distribution functions (CDFs) corresponding to independent univariate priors:
$\psi_{q,p}(u_{p})=F^{-1}(\pi_{q,p})(u_{p}).$ (4)
Here, $F^{-1}(\pi_{q,p}):[0,1]\to\mathcal{Z}$ is the generalized inverse CDF associated with a parameterized latent distribution, $\pi_{q,p}$. This is valid under standard regularity assumptions on $F$, since it is a measurable, monotone, increasing function, and thus $F^{-1}(u)=\inf\left\{z\in\mathcal{Z}:F(z)\geq u\right\}$ is a measurable mapping between the two probability spaces.
We parameterize the latent distribution $\pi_{q,p}$ via exponential tilting of a base prior $\pi_{0}$:
$\pi_{q,p}(z)=\frac{\exp(f_{q,p}(z))}{Z_{q,p}}\pi_{0}(z),$ (5)
where $f_{q,p}(z)$ is a learned energy function and $Z_{q,p}$ is the finite normalizing constant (partition function) that ensures $\pi_{q,p}$ defines a valid probability. The inspiration for this exponential tilting form can be traced to the energy-based prior introduced by Pang et al. (2020), adapted to the univariate case. The base prior $\pi_{0}$ serves as a reference measure that guides the form of the learned distribution $\pi_{q,p}$, allowing for flexible reweighting via the energy function $f_{q,p}$.
For $\pi_{q,p}$ to define a valid probability distribution, it must satisfy countable additivity while being non-negative and normalized. Assuming $f_{q,p}:\mathcal{Z}\to\mathbb{R}$ is $\mathcal{B}(\mathcal{Z})$-measurable with respect to the reference density $\pi_{0}$, this requires exponential integrability, i.e.,
@@
$\text{such that}\quad Z_{q,p}$ $=\int_{\mathcal{Z}}\exp(f_{q,p}(z))\,d\pi_{0}(z)<\infty.$ (6)
In practice, $f_{q,p}$ can be parameterized in various ways, made possible by the recent development of Kolmogorov-Arnold Networks (KANs) (Liu et al., 2024). For instance, smoothing splines, such as those used by KANs or deep spline networks, (Bohra et al., 2020). Splines are typically defined on bounded domains, allowing the support of $\pi_{q,p}$ to be explicitly controlled by the user.
The partition function $Z_{q,p}$ is often intractable in energy-based models. However, in this univariate setting, where the energy function is smooth and defined on a bounded domain $\left[z_{\text{min}},z_{\text{max}}\right]$, it can be accurately approximated using Gauss-Kronrod quadrature, (Laurie, 1997):
@@
where $G(z)=\exp(f_{q,p}(z))\pi_{0}(z)$, and $\{z_{i}^{\text{node}},w_{i}\}_{i=1}^{N_{\text{quad}}}$ are the quadrature nodes and weights for the interval $\left[z_{\text{min}},z_{\text{max}}\right]$. The smoothness of $G$, together with its bounded support, makes this approximation highly accurate.
With a valid form for the normalized CDF, inverse transform sampling (ITS) can be used to draw samples from $\pi_{q,p}$ as deterministic transformations of uniform noise. This preserves the deterministic structure of KART while enabling flexible, learnable latent distributions. Inverse transform sampling remains effective even for functions that are not Riemann integrable, provided they are integrable in the Lebesgue sense, typical for the functions in KAN literature.
As a result, the model retains a coherent probabilistic interpretation of KART that is consistent with the energy-based framework proposed by Pang et al. (2020), which is well suited to maximum likelihood training. However, this approach is significantly limited, as it ignores inter-dimensional dependencies in the latent space and results in an axis-aligned joint distribution. We address this inflexibility in Sec. 2.3.2 by relaxing our adherence to KART.

### v7

@@ 1 Introduction
Deep generative modeling is a core subfield of machine learning that enables systems to learn complex data distributions, generate novel data, and support a wide range of tasks including simulation, compression, and representation learning. Two prominent and widely studied approaches are Variational Autoencoders (VAEs) (kingma2022autoencodingvariationalbayes) and Generative Adversarial Networks (GANs) (goodfellow2014generativeadversarialnetworks). Both frameworks assume that each data point is generated from a low-dimensional latent vector via a top-down neural network, where the latent vector is typically drawn from a simple, non-informative prior, such as a uniform or isotropic Gaussian distribution.
@@
Latent energy-based prior models (EBMs) (pang2020learninglatentspaceenergybased) extend this framework by learning an expressive, data-dependent prior directly in the latent space. This prior can be trained jointly with the top-down generator using maximum likelihood estimation (MLE), where the EBM learns by the statistical discrepancy between samples from the prior and posterior distributions, following a paradigm similar to Empirical Bayes. Each learning iteration requires Markov Chain Monte Carlo (MCMC) sampling from both the prior and posterior, specifically the Unadjusted Langevin Algorithm (ULA) (Brooks2011; Rossky1978; Roberts2002), which often offers greater modeling flexibility compared to the amortized inference methods of VAEs.
@@
Despite being extendable to several high-impact research directions, such as multimodal generative modeling (yuan2024learningmultimodallatentgenerative), and interpretable classification, text-to-text, and meta-learning (pang2021latentspaceenergybasedmodeltext), (which involves rapid adaptability and few-shot generalization to out-of-distribution tasks), latent EBMs face critical limitations due to their reliance on MCMC sampling.
@@
Firstly, MCMC is inherently iterative, and algorithms such as ULA require repeated gradient passes through the model, which introduces significant computational overhead during both training and inference. Secondly, ULA is best suited for sampling from smooth, unimodal distributions and can struggle to explore the multimodal latent distributions that are typical in latent variable generative modeling. Moreover, the framework presented by pang2020learninglatentspaceenergybased requires separate ULA samplers for the prior and posterior. This introduces additional experimental complexity since ULA is sensitive to hyperparameter tuning: step sizes that are too small lead to poor mixing and convergence, while larger step sizes can accumulate discretization bias due to its use of the Euler-Maruyama method (Wilkie_2004) to evolve the underlying Langevin equation.
@@
Additionally, it is unclear how the unique separation between the prior and generator networks can be leveraged for interpretability or training efficiency. Unlike the framework introduced by Karniadakis2021, which describes how scientific inductive and learning biases can be embedded into neural networks to reduce their reliance on data and improve training efficiency, generalization, and explainability, (all of which are critical for ensuring trust in Artificial Intelligence), it remains unclear how the EBM prior can be effectively aligned with the posterior to achieve similar effects. Although ideas from mechanistic interpretability, as reviewed by bereska2024mechanisticinterpretabilityaisafety, could potentially be applied, there is currently no established method to leverage the unique strengths of the EBM prior framework for this purpose.
@@ 1.1 Inductive biases and fast sampling
To address these limitations, we reimagine the EBM prior into a novel univariate or element-wise model, drawing inspiration from Kolmogorov-Arnold Networks (KANs) (liu2024kankolmogorovarnoldnetworks) and deep spline networks 9264754, which replace the sigmoidal activations used in multilayer perceptrons MURTAGH1991183 with spline interpolations. By representing the latent space with univariate functions and distributions, we enable fast, (single forward pass), and exact prior sampling with the inverse transform method (devroye1986nonuniform) and improve interpretability for domain experts who may not be deeply familiar with machine learning. This allows the EBM prior to be aligned with the posterior before training begins, and for latent distributions to be plotted, viewed, and potentially understood with the symbolic regression approach proposed by liu2024kankolmogorovarnoldnetworks.
In Sec. 3.1, we show that simple datasets can allow for efficient training with importance sampling (IS) (goertzel1949quota). This is a technique for estimating posterior expectations that is otherwise avoided in the literature due to its high variance under severe prior-posterior mismatch and the curse of dimensionality, (dupuis2004importance), which can be avoided in our model due to the low dimensionality of the latent space. Moreover, we show that the trained priors are easy to visualize and recover, offering new opportunities for latent discovery and knowledge transfer to models with different latent structures.
To derive our new model, which we call the Kolmogorov-Arnold Energy Model (KAEM), we first introduce a novel adaptation of the Kolmogorov-Arnold Representation Theorem (KART) (Arnold2009) by reinterpreting its inner functions as Markov kernels between probability spaces. Given its finite and definitive nature, KART can provide structural bias and reduce the reliance of architecture design on intuition and automated tuning.
@@ 1.2 Population-based Langevin sampling and Thermodynamic training
However, while ITS enables fast inference, posterior sampling must still be conducted using ULA when the IS estimator is rendered ineffective due to variance. Prior work addressing poor MCMC convergence and multimodal exploration in EBMs has typically relied on diffusion models, such as the hierarchical EBM by cui2024learninglatentspacehierarchical and the diffusion-assisted EBM by zhang2023persistentlytraineddiffusionassistedenergybased.
To maintain prior interpretability and fast inference post-training, we instead introduce posterior annealing as a novel and alternative strategy. This involves decomposing the posterior into a sequence of power posteriors, presented by 2d186535-310b-3428-8c6f-e4ef298483d7, and recasting ULA with population-based MCMC methods (PT; Marinari1992; Hukushima1996).
The training criterion is derived from the discretized Thermodynamic Integral and the Steppingstone Estimator (SE), as presented by CALDERHEAD20094028; Annis2019 to ensure the impact of annealing on learning dynamics and parameter updates. Although this might be less expressive than hierarchical EBMs, it ensures that inductive biases remain easy to embed into the latent prior and interpret after training; and unlike denoising, which scales sequentially, annealing can scale by adding more annealed distributions in parallel.
While ULA is computationally efficient and enables more flexible chain dynamics, it does not preserve detailed balance or reversibility. This can be recovered by incorporating Metropolis-Hastings (MH) adjustments (13ab5b5e-0237-33fb-a7a8-6f6e4e0d4e0f) via the Metropolis-adjusted Langevin Algorithm (MALA) (10.1111/j.1467-9868.2010.00765.x). However, ULA remains the preferred method in our work, motivated by its efficiency and effectiveness, demonstrated by pang2020learninglatentspaceenergybased and further supported by the findings of nijkamp2019learningnonconvergentnonpersistentshortrun; nijkamp2019anatomymcmcbasedmaximumlikelihood, with regards to how short-run, biased MCMC methods can still be highly effective for sampling from EBMs.
@@
1.3 Scope of experiments
KAEM was compiled and differentiated using Julia’s Reactant (Reactantjl2025) and Enzyme (moses2020enzyme; moses2021enzyme_gpu; moses2022enzyme_parallel) packages, which optimizes and lowers Julia code to Multi-Level Intermediate Representation (MLIR) (lattner2020mlircompilerinfrastructureend), thus offering best-in-class performance. However, Gaussian Radial Basis Functions (RBF KANs) (li2024kolmogorovarnoldnetworksradialbasis), were adopted instead of the cubic B-spline bases proposed by liu2024kankolmogorovarnoldnetworks, as RBFs are Reactant-compatible and more efficient for GPU implementation. They have also been proven to provide an optimal solution to the regularized approximation/interpolation problem by 58326.
The experiments in Sec. 3.2 demonstrate that our mixture prior model is able to generate coherent, (although low-quality), samples after a few training epochs on RGB images, while still supporting efficient, exact sampling via component-wise ITS. In contrast, annealing incurred substantial computational overhead, but produced small improvements in sample quality, rendering its recommendation context-dependent. However, this may change as a larger number of power posteriors, longer training durations, and more advanced posterior sampling algorithms become practical with the hardware advances proposed by exa2024. Their architecture enables fine-grained control over data flow and parallelized element-wise nonlinearity which is naturally well suited to supporting KAEM at scale. Their XPUs also provide function/layer fusion capabilities that may improve the potential of univariate priors and adherence to KART.
@@ 2 The Kolmogorov-Arnold Energy Model
The Kolmogorov-Arnold Representation theorem (KART) (Arnold2009) forms the basis of KAEM. A structured proof of the theorem has been provided by dzhenzher2022structuredproofkolmogorovssuperposition.
@@ 2.1 Univariate energy-based prior
The functions $\psi_{q,p}$ can be conveniently translated to the framework of inverse transform sampling (ITS) (devroye1986nonuniform). We interpret $\psi_{q,p}$ as the application of an inverse cumulative density function (CDF) to a uniform random variable to recover a sample from a univariate latent density, $p_{q,p}$:
$z_{q,p}=\psi_{q,p}(u_{p})=F^{-1}_{\pi_{q,p}}(u_{p})\quad\Leftrightarrow\quad z_{q,p}\sim p_{q,p}(z).$ (2)
Here, $F^{-1}_{\pi_{q,p}}:[0,1]\to\mathcal{Z}_{q,p}$ is the generalized inverse CDF associated with a parameterized latent measure, $\pi_{q,p}$ in the latent space, $\mathcal{Z}_{q,p}\subset\mathbb{R}$. We have therefore interpreted $u_{p}\sim\text{Unif}(0,1)$ in Eq. 1 as a uniform random variable in the probability space, $\left([0,1],\mathcal{B}([0,1]),\mu\right)$, where $\mathcal{B}\left([0,1]\right)$ is the Borel $\sigma$-algebra on $[0,1]$ and $\mu$ is the uniform probability measure on $[0,1]$.
This is valid under standard regularity assumptions on $F$, since it is a measurable, monotone, increasing function, and thus $F_{\pi_{q,p}}^{-1}(u)=\inf\left\{z\in\mathcal{Z}_{q,p}:F_{\pi_{q,p}}(z)\geq u\right\}$ is a measurable pushforward (or Markov kernel) between the two probability spaces:
$\left([0,1],\mathcal{B}\left([0,1]\right),\mu\right)\;\underset{F_{\pi_{q,p}}^{-1}}{\longrightarrow}\;\left(\mathcal{Z}_{q,p},\mathcal{B}\left(\mathcal{Z}_{q,p}\right),\pi_{q,p}\right),$ (3)
We parameterize the latent measure $\pi_{q,p}$ via exponential tilting of a base measure $\pi_{0}$:
$d\pi_{q,p}(z)=\frac{\exp(f_{q,p}(z))}{Z_{q,p}}d\pi_{0}(z),$ (4)
where $f_{q,p}(z)$ is a learned energy function and $Z_{q,p}$ is the finite normalizing constant (partition function) that ensures $\pi_{q,p}$ induces a valid probability density, $p_{q,p}$, for each latent feature.
The inspiration for this exponential tilting form can be traced to the energy-based prior introduced by pang2020learninglatentspaceenergybased, adapted to the univariate case. The base prior $\pi_{0}$ serves as a reference that guides the form of the learned $\pi_{q,p}$, allowing for flexible reweighting via the energy function $f_{q,p}$. For $\pi_{q,p}$ to define a valid density, which we denote $p_{q,p}$, we require exponential integrability:
@@
$\text{such that}\quad Z_{q,p}$ $=\int_{\mathcal{Z}_{q,p}}\exp(f_{q,p}(z))\,\pi_{0}(dz)<\infty.$ (5)
In practice, $f_{q,p}$ can be parameterized in various ways, made possible by the recent development of Kolmogorov-Arnold Networks (KANs) (liu2024kankolmogorovarnoldnetworks). For instance, smoothing splines, such as those used by KANs or deep spline networks, (9264754). Splines are typically defined on bounded domains, allowing the support of $\pi_{q,p}$ to be explicitly controlled by the user using the KAN grids. These grids must be periodically updated during training according to input activations, following the implementation described by liu2024kankolmogorovarnoldnetworks.
The partition function $Z_{q,p}$ is often intractable in energy-based models. However, in this univariate setting, where the energy function is smooth and defined on a bounded domain $\left[z_{\text{min}},z_{\text{max}}\right]$, it can be accurately approximated using Gauss-Kronrod quadrature, (gauss-quad):
@@
where $G(z)=\exp(f_{q,p}(z))\pi_{0}(z)$, and $\{z_{i}^{\text{node}},w_{i}\}_{i=1}^{N_{\text{quad}}}$ are the quadrature nodes and weights for the interval $\left[z_{\text{min}},z_{\text{max}}\right]$. The availability of $Z_{q,p}$ enables ITS, and the smoothness of $G$, together with its bounded support, makes this approximation accurate.
With this valid form for the normalized CDF, inverse transform sampling (ITS) can be used to draw samples from $p_{q,p}$ as measurable transformations of uniform noise. This preserves the deterministic structure of KART while enabling flexible, learnable latent distributions, well-founded so long as the KAN functions admit exponential integrability.
As a result, the model retains a coherent probabilistic interpretation of KART that is consistent with the energy-based framework proposed by pang2020learninglatentspaceenergybased, which is well suited to maximum likelihood training. However, this approach is significantly limited, as it ignores inter-dimensional dependencies in the latent space and results in an axis-aligned joint distribution. We address this inflexibility in Sec. 2.3.2 by relaxing our adherence to KART.
2.1.1 Multivariate prior notation
From this point forward, we formulate our derivations over the multivariate probability space, $\bigl(\mathcal{Z},\mathcal{B}(\mathcal{Z}),\mathbb{P}_{\bm{\theta}}\bigr),$ where $\mathbb{P}_{\bm{\theta}}$ denotes a product measure parameterized by abstract model parameters, $\theta$, and the latent space is defined as a Cartesian product:
$\mathbb{P}_{\bm{\theta}}\;:=\;\bigotimes_{q=1}^{2n_{z}+1}\;\bigotimes_{p=1}^{n_{z}}\pi_{q,p},\quad\quad\quad\mathcal{Z}\;=\;\prod_{q=1}^{2n_{z}+1}\prod_{p=1}^{n_{z}}\mathcal{Z}_{j,k}\;\subset\;\mathbb{R}^{(2n_{z}+1)\times n_{z}},$ (7)
Each $\pi_{q,p}$ is a univariate probability measure on $\mathcal{Z}_{q,p}$ defined via exponential tilting, as introduced in Eq. 4. Under $\mathbb{P}_{\bm{\theta}}$, the latent variables $\{z_{q,p}\}_{q,p}$ are mutually independent, and the full latent variable
$\bm{z}\;=\;\bigl(z_{q,p}\bigr)_{q=1,\ldots,2n_{z}+1;\;p=1,\ldots,n_{z}}$
is a matrix-valued random variable whose entries correspond to independent univariate features.